### v1

# Exploring System 1 and 2 communication for latent reasoning in LLMs

Abstract
Should LLM reasoning live in a separate module, or within a single model’s forward pass and representational space? We study dual-architecture latent reasoning, where a fluent Base exchanges latent messages with a Coprocessor, and test two hypotheses aimed at improving latent communication over Liu et al. (2024b): (H1) increase channel capacity; (H2) learn communication via joint finetuning. Under matched latent-token budgets on GPT-2 and Qwen-3, H2 is consistently strongest while H1 yields modest gains. A unified soft-embedding baseline—a single model with the same forward pass and shared representations, using the same latent-token budget—nearly matches H2 and surpasses H1, suggesting current dual designs mostly add compute rather than qualitatively improving reasoning. Across GSM8K, ProsQA, and a Countdown stress test with increasing branching factor, scaling the latent-token budget beyond small values fails to improve robustness. Latent analyses show overlapping subspaces with limited specialization, consistent with weak reasoning gains. We conclude dual-model latent reasoning remains promising in principle, but likely requires objectives and communication mechanisms that explicitly shape latent spaces for algorithmic planning.
[…]
Julian Coda-Forno at
1 Introduction
Large language models (LLMs) trained with web-scale pretraining and alignment have achieved impressive zero-shot reasoning capabilities across diverse tasks (Dubey et al., 2024; Hurst et al., 2024; Yang et al., 2025; Liu et al., 2024a). Despite their strong performance, LLMs are often seen as “fast and fluent” rather than genuinely deliberative — resembling the intuitive, System-1 side of dual-process theories of cognition (Kahneman, 2011). Indeed, recent surveys explicitly describe progress in reasoning LLMs as a shift from System-1-like heuristics to System-2-style deliberation (Li et al., 2025b), highlighting the need for architectures that support more structured reasoning.
The dominant approach today is chain-of-thought (CoT) reasoning (Wei et al., 2022; Liu et al., 2024a), where intermediate steps are verbalized in natural language. While effective, CoT incurs substantial token-level overhead, limits abstraction bandwidth, and constrains inference unnecessarily to the sequential, symbolic space of text (Chen et al., 2025; Qu et al., 2025). As model sizes and context window grow, these inefficiencies become increasingly prohibitive, motivating the search for more compact and expressive reasoning representations.
Latent reasoning (Hao et al., 2024; Liu et al., 2024b; Geiping et al., 2025) offers an alternative that enables the model to perform multi-step inference internally within its continuous hidden states, surfacing only the final answer. This paradigm promises two key advantages. First, reasoning in high-dimensional embeddings provides vastly greater expressive bandwidth than token sequences, potentially allowing richer intermediate computations (Zhang et al., 2025). Second, for combinatorial problems, operating over structured latent abstractions can dramatically reduce the effective search space (Geiping et al., 2025). Such ideas find parallels in cognitive science, where humans are believed to reason in internal “mentalese" before translating thoughts into language (Fodor, 1975).
Most latent-reasoning methods still ask a single network to do both fast association and slow deliberation; e.g., Coconut (Hao et al., 2024) feeds the model’s last hidden state back as input, forcing the same space to support next-token prediction and a putative “language of thought,” creating a representational tug-of-war. Neuroscience evidence instead points to partially distinct substrates (PFC for deliberative control; striatal circuits for habitual responses (Miller and Cohen, 2001; O’Reilly and Frank, 2006; Dolan and Dayan, 2013)), suggesting that separating roles could help. KV-cache Coprocessors (Liu et al., 2024b) fit this separation, but reported gains are limited; we argue the bottleneck is latent communication between modules.
We therefore revisit the KV-Coprocessor design with two changes aimed at strengthening communication: (i) frozen-Base cache augmentation, where the Coprocessor writes cache edits that reach all layers of the Base, and (ii) co-finetuning, where the Base and Coprocessor are trained jointly to make the Base “listen” to latent messages. With matched token budgets and latent counts $N_{L}$, we evaluate two model families on pretraining and reasoning (GSM8K, ProsQA, Countdown) and analyze whether latents specialize. This lets us test whether these architectures deliver genuine “latent reasoning” rather than merely adding compute.
Overall, we contribute to the literature in the following ways:
  1. 1.
    We introduce two communication-oriented variants of Liu et al. (2024b)’s dual architecture and show better performance under matched latent-token budgets.
  2. 2.
    Using reasoning and latent-space analyses, we find that gains in current dual architectures largely reflect added capacity rather than structured reasoning; latents show limited specialization, suggesting that objectives encouraging specialization may be necessary for latent reasoning.
2 Related work
Reasoning in latent space. Latent-space methods shift multi-step inference from tokenized CoT into a small set of latent variables, decoupling compute from text length. Representative approaches include Coconut, which feeds continuous “thoughts” back into the model (Hao et al., 2024); differentiable KV-cache augmentation, which separates a Base token predictor from a Coprocessor that edits the cache (Liu et al., 2024b); and compressed/implicit CoT that learns dense contemplation tokens or plans without emitting text (Cheng et al., 2024; Kurtz et al., 2024). Recent surveys synthesize this trend and its claimed benefits (Li et al., 2025a; Zhu et al., 2025). Despite clear token-efficiency gains, reported improvements on reasoning are mixed, motivating our study of how to train dual-model systems for effective latent communication.
Communication and coordination between models. Our focus on “latent communication” between a Base and a Coprocessor connects to two strands. First, teacher–student transfer suggests that sharing an initialization can facilitate implicit trait transfer, even without explicit supervision (Cloud et al., 2025); this supports initializing the Coprocessor from the Base to align internal representations. Second, multi-agent prompting frameworks use inter-model dialogue to improve reliability (Du et al., 2023; Liang et al., 2023), though recent analyses find gains can be brittle or dataset-dependent (Wang et al., 2024). Finally, our unified soft-embedding baseline is grounded in continuous-prompt methods that endow a single network with extra latent capacity via trainable prefix/prompt vectors (Lester et al., 2021; Li and Liang, 2021; Goyal et al., 2023), providing a strong, parameter-efficient alternative to dual-model designs.
3 Methods
3.1 Problem setting
Figure 1: Overview of the deliberation–in–KV-cache architecture of Liu et al. (2024b) and our two variants designed to strengthen cross-module communication.
Setup and notation. Let $B_{\theta}$ be a frozen, pretrained LLM (Base) with parameters $\theta$. Given a prompt $x$ and target $y$, a forward pass of $B_{\theta}$ produces per-layer key–value caches $\{(K_{\ell}(x;\theta),V_{\ell}(x;\theta))\}_{\ell=1}^{L}$, abbreviated $KV_{\theta}(x)$. A Coprocessor $C_{\phi}$ with parameters $\phi$ reads $KV_{\theta}(x)$ together with $N_{L}$ learnable soft tokens and emits a latent sequence $Z\in\mathbb{R}^{N_{L}\times d}$. At decoding time, $Z$ is injected back into $B_{\theta}$ (in a manner that depends on the variant below), after which the Base generates $y$.
Objective. The training goal is to learn latents that improve conditional likelihood:
$\max_{\phi\;\;(\text{and possibly }\theta)}\;\mathbb{E}_{(x,y)\sim\mathcal{D}}\big[\log p_{\theta}\big(y\mid x,\operatorname{inject}(Z;KV_{\theta}(x))\big)\big],$
where $Z=C_{\phi}\big(KV_{\theta}(x),N_{L}\text{-soft}\big)$ and $\operatorname{inject}(\cdot)$ denotes the chosen mechanism for feeding $Z$ (or the Coprocessor’s cache) back to the Base.
Pipeline (as in Liu et al. (2024b)). We follow the standard three-stage process (cf. Fig. 1, left):
  1. (i)
    KV-cache generation. Run $B_{\theta}$ on $x$ once to obtain $KV_{\theta}(x)$.
  2. (ii)
    Latent augmentation. Run $C_{\phi}$ on $KV_{\theta}(x)$ plus $N_{L}$ learnable soft tokens to produce a latent sequence $Z$. In Liu et al. (2024b), only the final layer’s hidden states of $C_{\phi}$ are used as input embeddings to $B_{\theta}$.
  3. (iii)
    Decoding. Inject the augmentation ($KV_{\theta}(x)$ and $Z$) into $B_{\theta}$ and decode $y$. Unless otherwise noted, only $\phi$ and the soft-token embeddings receive gradients.
3.2 Experimental variants
We test two hypotheses aimed at strengthening communication between modules and isolating where the gains arise. Each hypothesis states a falsifiable prediction at matched latent-token budgets.
Hypothesis 1 — Frozen-Base KV augmentation. Change: Instead of converting the Coprocessor’s last layer’s output into input embeddings, we concatenate its per-layer cache to the Base cache at injection time. Let $KV_{\phi}(x)$ denote the Coprocessor’s cache produced from $KV_{\theta}(x)$ and the $N_{L}$ soft tokens. Decoding uses
$K_{\ell}(x;\theta)\,;\,K_{\ell}(x;\phi)]\quad\text{and}\quad[V_{\ell}(x;\theta)\,;\,V_{\ell}(x;\phi)]\qquad\forall\,\ell=1,\dots,L,$
i.e., concatenation along the sequence (cache) dimension. [^1] We denote sequence-axis concatenation by $\cdot\,$$\cdot$]. Optimization: $\theta$ remains frozen; only $\phi$ and the soft tokens are trained. Motivation: With a frozen Base, steering only via input embeddings gives the Coprocessor influence at the first hidden layer. Cache-level augmentation propagates the latent signal through all layers, potentially enabling richer, layer-wise “latent communication”. Prediction: With $\theta$ frozen, cache augmentation will outperform embedding-only feedback (Liu et al., 2024b) at matched latent-token budgets $N_{L}$.
Hypothesis 2 — Co-finetuned dual-model. Change: Same injection mechanism as the reference system in Liu et al. (2024b) (latents as input embeddings to the Base), but we unfreeze the Base. Optimization: Update both $\theta$ and $\phi$ jointly (plus soft tokens) under the log-likelihood objective above. Motivation: Joint training allows $B_{\theta}$ to learn to “listen” to the Coprocessor’s messages rather than treating them as exogenous noise. Although this increases the number of trainable parameters, it isolates whether improved cross-module interaction—rather than mere extra compute—drives downstream gains. Prediction: For the same latent injection as Liu et al. (2024b), co-finetuning will outperform the frozen-Base counterpart at matched $N_{L}$, reflecting learned communication rather than added compute.
Implementation note. Both variants are realized within the same three-pass schedule used throughout the paper (§4.1): Base pass to form $KV_{\theta}(x)$ $\rightarrow$ Coprocessor pass to form $Z$ (and, for Hyp. 1, $KV_{\phi}(x)$) $\rightarrow$ Base decode with injection. This preserves a clean separation between next-token prediction and latent computation.
4 Experimental analysis:
4.1 Large scale data training
We replicate the deliberation KV-cache pipeline of Liu et al. (2024b), but adapt it to a tighter compute budget and to two model families.
Base models: GPT-2 (124 M) (Radford et al., 2019) and Qwen-3 (0.6 B) (Yang et al., 2025) replace the 2-B-parameter Gemma used in the Liu et al. (2024b) paper. This choice allows us to complete all runs within a 3-day window on an 8 × A100 (80 GB) node.
Token budget: Liu et al. train for $\approx 2\times 10^{11}$ tokens (sequence 2048, batch 1024, 100 k steps). We scale this down to 40 B tokens for GPT-2 and 8 B tokens for Qwen-3, which we found sufficient to reproduce their qualitative trends without exceeding our hardware envelope.
Dataset. All large-scale training uses FineWeb-Edu-100BT (Lozhkov et al., 2024).
Sequence & latent parameters: We use sequence length S=1024S{=}1024, latent augmentations per sequence M=64M{=}64, ahead tokens $N_{A}{=}16$ for back-propagating the loss into the Coprocessor (Liu et al. (2024b) use $2048/128/16$), and $N_{L}{=}16$ latents. Figure 2 C/D show that validation perplexity decreases almost linearly as $N_{L}$ grows under Hypothesis 2. Training cost scales with the effective per-example context $S+M\!\cdot\!(N_{L}{+}N_{A})$, so larger $N_{L}$ substantially raises wall-clock. Within our budget, $N_{L}{=}16$ provided a practical operating point, while Liu et al. (2024b) report their strongest results around $N_{L}{=}32$ on larger bases.
Three-pass training loop implementation (Fig. 5): Liu et al. (2024b) describe “an efficient training framework … in one forward pass”, enabled by a custom attention mask that scales to large datasets. Such a mask must support multiple augmentations (M) per sequence; otherwise only one augmentation is trained, yielding two orders of magnitude less signal per token. However, a single forward pass could allow the Base model to shortcut the intended latent computation—performing both next-token prediction and latent augmentation itself, thereby collapsing the two representational spaces and defeating the purpose of the technique. We cannot verify whether Liu et al. (2024b) avoid this, as their code is not public; the attention masks they describe imply that only one model processes the pass, but the exact implementation remains unclear. To keep our dataset large while preventing this shortcut, we adopt a strict three-pass loop (Figure 5) whose attention masks still allow M augmentations of each sequence.
Only the Coprocessor (and, in Hypothesis 2, the Base model) receives gradients. Although this three-pass schedule incurs a 3x wall-clock penalty compared with the unknown “single-pass” implementation, it preserves a clean separation between next-token prediction and latent reasoning—the property we wish to evaluate.
After convergence, models are evaluated greedily on standard pretraining benchmark for small LLMs: HellaSwag, ARC-Easy, SocialIQA, PIQA, and Winogrande.
4.1.1 Baselines
Baseline 1 - Base model + continued pretraining: Liu et al. (2024b), compare their dual-model to the initial base checkpoint, meaning the dual system has seen far more data. Instead, we keep the architecture unchanged and simply continue pretraining the base model on the same number of tokens used by our dual-model runs.
Baseline 2 - Base model + $N_{L}$ soft embeddings: To test whether a single network can absorb the “System-2” role, we continue pretraining the base model alone while attaching the same number of new $N_{L}$ learnable soft tokens used by the Coprocessor setup (similar in spirit to Goyal et al. (2023), but with untied, slot-specific embeddings rather than a single shared token). This contrasts a unified versus dual-network design under identical data and latent embedding budgets. See App. 6 for visualization.
Note. Neither baseline is strictly parameter-matched—our dual-model has roughly twice as many weights—yet they provide useful sanity checks on data scaling and on the value of a separate Coprocessor.
4.1.2 Results
A                    B
C                    D
Figure 2: Validation perplexity whilst training on the FinWeb-Edu-100BT corpus. A: GPT-2 variants (using $N_{L}{=}16$ where applicable). B: Qwen-3 variants (using $N_{L}{=}16$ where applicable). C: Ablating number of latents $N_{L}$ of the GPT-2 Coprocessor for Hypothesis 2. D: Ablating number of latents $N_{L}$ of the Qwen-3 Coprocessor for Hypothesis 2.
Model Hellaswag ARC-Easy Social IQA PIQA Winogrande
GPT-2 variants
[…]
GPT-2 + soft embeddings 30.7 (+0.0) 55.6 (+1.4) 37.7 (–0.1) 64.6 (+0.1) 51.9 (+1.4)
Liu et al. (2024b) 29.0 (–1.7) 45.0 (–9.2) 37.4 (–0.4) 62.1 (–2.4) 50.7 (+0.2)
Hypothesis 1 29.4 (–1.3) 48.2 (–6.0) 38.7 (+0.9) 62.9 (–1.6) 52.3 (+1.8)
Hypothesis 2 31.2 (+0.5) 55.8 (+1.6) 38.2 (+0.4) 65.3 (+0.8) 52.1 (+1.6)
Qwen variants
[…]
Qwen + soft embeddings 25.9 (–0.7) 37.2 (+2.7) 34.6 (–0.7) 55.3 (+0.2) 50.4 (–1.6)
Liu et al. (2024b) 31.8 (+5.2) 36.5 (+2.0) 35.4 (+0.1) 58.1 (+3.0) 52.0 (+0.0)
Hypothesis 1 35.2 (+8.6) 45.0 (+10.5) 39.0 (+3.7) 61.0 (+5.9) 55.7 (+3.7)
Hypothesis 2 37.3 (+10.7) 53.1 (+18.6) 41.9 (+6.6) 63.4 (+8.3) 51.1 (–0.9)
Table 1: Model performance on standard small LLM benchmarks and $\Delta$ with continued pretraining baselines.
Qualitative replication of Liu et al. (2024b). Substituting the 2-B parameter Gemma with much smaller GPT-2 (124 M) and Qwen-3 (0.6 B) still produces the qualitative trends reported by Liu et al. (2024b)(Fig. 2A/B). Perplexity (ppl) drops by $\simeq$ 2 for GPT-2 and $\simeq$ 2.5 for Qwen-3—an order-of-magnitude larger reduction than that reported in the original paper setup. On GPT-2 the dual-model also raises mean benchmark accuracy by +2.0 percentage point ($p$) (Table 1). For Qwen-3, however, lower perplexity does not translate into higher benchmark scores; continued pretraining alone even degrades accuracy. Given Qwen-3’s heavy post-training, such brittleness under distribution shift is expected. We therefore report all numbers relative to the continued-pretraining baseline.
Both hypotheses outperform Liu et al. (2024b). Relative to Liu et al.’s dual-model, Hypothesis 1 (frozen Base, cache concatenation) improves average accuracy by +1.5 $p$ on GPT-2 and +4.4 $p$ on Qwen-3. Hypothesis 2 (co-finetuned) improves by +3.7 $p$ (GPT-2) and +6.6 $p$ (Qwen-3). In ppl, the variants are $\approx 2$ and $\approx 4$ lower, respectively (Fig. 2A/B).
Liu model versus data-matched continued pretraining. Relative to our data-matched Baseline 1 (same number of training tokens), the Liu et al.’s dual-model exhibits higher ppl ($\approx{+}1.5$) and mixed accuracy: 2.7-2.7 $p$ on GPT-2 but +2.1+2.1 $p$ on Qwen-3. Averaged across families the net change is $\approx{-}0.3$ $p$, underscoring the value of stricter baselines.
Dual models versus the “soft-embedding” baseline. Our second baseline—Base model + $N_{L}$ soft embeddings—keeps the architecture unified yet endows the network with the same latent budget. Against this control the dual-model results are underwhelming:
  • Perplexity: Hyp. 1 has higher ppl than the soft-embedding model for both families; Hyp. 2 lowers it only marginally.
  • GPT-2 benchmarks: Averaged over the five tasks in Table 1, Hyp. 1 trails by 1.8-1.8 $p$ (46.3 vs. 48.1%), while Hyp. 2 is only +0.4+0.4 $p$ higher (48.5%).
  • Qwen-3 benchmarks: Dual models fare better: Hyp. 1 +6.5+6.5 $p$ (47.2 vs. 40.7%); Hyp. 2 +8.7+8.7 $p$ (49.4%).
Summary and caveat. At first glance both dual-model variants look successful: they outperform the Liu et al. (2024b) baseline and, on Qwen-3, yield sizeable benchmark gains. But the picture shifts once we introduce the soft-embedding control. That unified model matches the dual systems in token budget and latent capacity while using only half as many trainable weights, yet it equals or exceeds Hyp. 1 and comes within a hair of Hyp. 2. This pattern indicates that the Coprocessor is not just “adding compute”—it is adding it inefficiently. A single LLM with the same aggregate parameter count would almost certainly do better. Accordingly, the next section turns to reasoning-specific benchmarks to see whether the dual architecture offers any benefit that a parameter-matched, soft-prompted model cannot already provide.
4.2 Reasoning evaluation: benchmarks and a controlled stress test
In this section we wanted to test if additional latent tokens improve reasoning or do they mostly add compute? We first compare our systems on GSM8K (Cobbe et al., 2021) and ProsQA (Hao et al., 2024), then use a controlled Countdown stress test (Pan et al., 2025) to probe robustness as combinatorial difficulty increases. We evaluate four systems: our reproduction of Liu et al., Hyp. 1 (frozen Base, cache-concat), Hyp. 2 (co-finetuned), and a single-model soft-embedding baseline with the same total latent budget $N_{L}$. For reference, we also compare against Coconut (GPT-2) (Hao et al., 2024). More implementation details can be found in App. 7.
Benchmarks. On GSM8K/ProsQA we follow the staged curriculum of Hao et al. (2024) and report accuracy after fine-tuning. Table 2 shows the best results ($N_{L}=12$ for this task). Figure 3A plots GSM8K accuracy versus total latents $N_{L}$; the corresponding ProsQA curves are deferred to App. 7, as this benchmark is near-saturated for competitive systems and adds limited insight.
Stress test: Countdown. Countdown lets us scale difficulty in a controlled, task-homogeneous way by increasing the operand count (branching factor grows rapidly with each operand). Unlike the benchmarks, we train without curriculum and sweep both $N_{L}\!\in\!\{1,2,4,8\}$ and operands $\in\{3,4,5\}$. An illustrative instance appears in the box below. Figure 3B plots accuracy vs. operands, combining GPT-2 and Qwen curves for clarity.
Countdown example with operands = 4 User: Using the numbers [19, 36, 55, 7][19,\,36,\,55,\,7], create an equation that equals 65. Assistant: Let me solve this step by step. <latent thinking>
<answer>
55+3671955+36-7-19 </answer>
Findings. (i) Rank order on GSM8K/ProsQA: Hyp. 2 $\geq$ soft-embedding $\gg$ Hyp. 1 $\approx$ Liu (Table 2). (ii) Scaling latents: Fig. 3A shows accuracy is largely flat as $N_{L}$ increases and can dip at larger $N_{L}$. This contrasts with Section 4.1, where perplexity decreased as $N_{L}$ grew, suggesting that extra latents help next-token prediction but do not translate into more reliable reasoning. Interestingly, Coconut (GPT-2) exhibits the same dipping trend, indicating the difficulty is not specific to dual-model designs. (iii) Stress test: In Fig. 3B, increasing operands sharply reduces accuracy for all systems; larger $N_{L}$ helps slightly up to $N_{L}=8$ but yields diminishing returns thereafter, and remains very close to the single-model soft-embedding baseline.
Table 2: Accuracy (%) on GSM8K and ProsQA after curriculum fine-tuning with $N_{L}=12$.
GPT-2 Qwen-3
Model GSM8K ProsQA GSM8K ProsQA
Coconut (B0) 34.1 97.0
Soft emb. (B1) 26.5 97.7 38.5 99.5
Liu et al. (B2) 16.5 52.6 22.4 81.0
Hyp. 1 (frozen, concat) 12.0 54.1 24.0 79.0
Hyp. 2 (co-finetuned) 31.5 99.0 38.6 99.5
Efficiency note (Hyp. 2 vs. Coconut). Coconut generates continuous thoughts sequentially with the same network; for a given input this requires approximately $N_{L}{+}1$ full forward passes (each time appending a new latent and re-processing), so compute and latency scale linearly with $N_{L}$ Hao et al. (2024). Our Hyp. 2 uses a strict three-pass schedule independent of $N_{L}$ (Base cache $\rightarrow$ Coprocessor latents $\rightarrow$ Base decode). Assuming similar model sizes, this yields an approximate forward-pass/FLOPs reduction of
$\text{speedup}\;\approx\;\frac{N_{L}+1}{3}\,$,
e.g., $\sim 5.7\times$ fewer full passes at $N_{L}{=}16$, while remaining batch-parallel (multiple augmentations per sequence can still be handled in one Coprocessor pass). In practice this removes Coconut’s serial bottleneck—while achieving comparable task accuracy on GPT-2 (GSM8K: 31.5 vs. 34.1; ProsQA: 99.0 vs. 97.0; Table 2, Fig. 3a)—enabling larger models and datasets. However, it is worth noting that Hyp. 2 also introduces roughly $2\times$ as many trainable parameters as Coconut.
A B
Refer to caption Refer to caption
Figure 3: Ablating the latent budget. A: GSM8K accuracy vs. total latents $N_{L}$ (GPT-2 and Qwen; Coconut shown for GPT-2). B: Countdown accuracy vs. operands (3, 4, 5) with lines for $N_{L}\!\in\!\{1,4,8,16\}$, merged across model families.
Summary. Across reasoning benchmarks and the controlled stress test, adding latents mostly buys FLOPs, not robust reasoning. A unified soft-embedding model closely matches the best dual-model at equal $N_{L}$, and further increasing $N_{L}$ rarely helps—sometimes hurts. This motivates the interpretability analysis in §4.3.
4.3 Interpretability: do latents specialize or collapse?
Reasoning can be viewed as composing distinct intermediate computations or modules (Andreas et al., 2016; Lake and Baroni, 2017; Veličković and Blundell, 2021). If our Hypothesis 2 Coprocessor truly supports such division of labor, different latents should occupy meaningfully different directions in representation space. If, instead, latents mostly scale confidence without new algorithmic structure, they will reuse the same span, yielding redundant computations (Olah et al., 2020; Elhage et al., 2022; Kornblith et al., 2019). We therefore analyze the Coprocessor’s last hidden layer (the signal fed back as input embeddings to the Base) and test whether latents specialize.
Diagnostic 1: Cross-capture heatmap. Intuition. If latents specialize, the variance of latent $j$’s activations should not lie in latent $i$’s subspace. Let $X_{i}\in\mathbb{R}^{n_{i}\times d}$ be row-centered activations for latent $i$ in this layer. For each $i$, we fit a minimal PCA subspace explaining at least $\tau=97\%$ of its variance and form the projector $P_{i}$. We then quantify how much of latent $j$’s variance falls into that subspace,
$H_{i,j}\;=\;\frac{\bigl\|X_{j}P_{i}\bigr\|_{F}^{2}}{\bigl\|X_{j}\bigr\|_{F}^{2}}\;\in[0,1],$
visualize $H$ as a heatmap (note $H_{i,i}\!\geq\!\tau$), and summarize with the mean off-diagonal capture
$\overline{H}_{\mathrm{off}}\;=\;\frac{1}{N_{L}(N_{L}-1)}\sum_{i\neq j}H_{i,j}\quad\text{(higher $\Rightarrow$ more redundancy; lower is better).}$
Diagnostic 2: silhouette (cluster separation by latent). We compute the silhouette score (Kaufman and Rousseeuw, 1990), $s\in[-1,1]$, which for each point compares how close it is to its own latent cluster versus the nearest other latent cluster; higher $s$ means tighter, better separated clusters. This complements diagnostic 1: Instead of looking at directional variance reuse across latents (orientation overlap), it looks at instance-level spatial separation/cohesion of latent-labeled clusters. We report the global average (formal definition and the mean per-latent scores in App. §8).
4.3.1 Results
Figure 4: Latent cross-subspace capture heatmaps (last Coprocessor layer). Each cell $(i,j)$ shows the fraction of variance of latent $j$ captured by the principal subspace of latent ii. A: Large-scale training. B: Fine-tuning on GSM8k with curriculum. C: Finetuning on countdown with operands = 4.
Large-scale training. Figure 4a is nearly uniform and bright off the diagonal, and the quantitative scores confirm this collapse-like behavior: mean off-diagonal capture $\overline{H}_{\text{off}}=0.9873$ (higher $\Rightarrow$ more redundancy), global silhouette s=0.1694s=-0.1694 with per-latent silhouettes mostly negative ($\{-0.26,\dots,-0.07\}$; full vector in Appendix 9). Together these indicate that occurrences labeled by different latents largely occupy the same subspaces and are not cluster-separated.
GSM8K with curriculum. Figure 4b shows the curriculum–tuned model; the no–curriculum variant is in Appendix 9. Relative to the large-scale pretraining task, fine-tuning reduces directional redundancy among latents: $\overline{H}_{\text{off}}$ drops to $0.962$ without curriculum and to $0.914$ with curriculum. However, cluster separation does not emerge: the global silhouette remains near zero in both task-tuned settings ($s\approx-0.031$). In short, curriculum nudges representations in the right direction—latents become less overlapping in orientation—but they still occupy essentially the same region of space (decorrelated directions without spatial separation).
Countdown (operands = 4). Countdown induces substantially better cluster separation (global silhouette $s=0.4531$), yet the cross-subspace redundancy remains high with $\overline{H}_{\text{off}}=0.9382$. This realizes the “separated but redundant spans” regime: clusters are distinct in Euclidean space (high silhouette) but their principal subspaces still explain most of each other’s variance (high off-diagonal capture). Interestingly, Fig 4c shows distinctiveness degrading beyond latent 8, aligning with Fig 3B where accuracy improves from $N_{L}\in\{1,4,8\}$ but drops at $N_{L}{=}16$.
Across all three settings the latents tend toward redundant computations: their subspaces are highly overlapping (high $\overline{H}_{\text{off}}$), even when the data task encourages class separation (Countdown). Task-aligned fine-tuning helps (curriculum << pretraining in $\overline{H}_{\text{off}}$), but not enough to yield “separated and specialized” latents (high silhouette and low off-diagonals).
5 Discussion
Building on Liu et al. (2024b), we cast the coprocessor architecture as a principled attempt to disentangle “System-1” token prediction from “System-2” abstract reasoning. Treating latent embeddings as a private communication channel between the two modules clarifies the conceptual link to dual-process theories in cognitive science and motivates our two new training hypotheses. However, in our setting, simply providing the channel (as in Liu et al. (2024b)) or strengthening it (our H1/H2) did not induce System-2-like computation.
Replacing last-layer hidden-state injection with full KV-cache concatenation yields modest perplexity and benchmark gains over the original design, yet falls short once stricter baselines are introduced. Its benefits appear sensitive to model family and vanish on harder reasoning tasks. Jointly updating both models produces the strongest results across all experiments, confirming that bidirectional adaptation facilitates cross-model communication. However, a unified network trained with the same soft-token budget narrows—sometimes erases—the gap, suggesting that current dual-system instantiations add compute inefficiently rather than unlocking qualitatively new reasoning abilities.
Our stricter baselines highlight scenarios where dual-model gains either disappear or can be matched closely by simpler means. The Countdown experiments further show that scaling the latent budget beyond eight tokens fails to deliver systematic robustness as combinatorial complexity explodes. Our interpretability analysis seem to corroborate this by showing that learned latents largely occupy overlapping representational subspaces: extra latents mostly amplify confidence rather than add new algorithmic structure.
Our negative results do not invalidate the dual-system framework; they indicate that how the two models exchange information remains an open problem, and that current instantiations do not create conditions for System-2-like computation to emerge. Promising directions include (i) designing objectives that explicitly reward diversity or orthogonality in latent representations to encourage broader search, and (ii) developing training schedules that preserve large-scale language competence while gradually shaping latent spaces for multi-step reasoning.